OpenStudy (anonymous):

Logarithms Evaluate the following expressions:

OpenStudy (anonymous):

\[\log _{2} 0.25\]

OpenStudy (anonymous):

\[\log _{10}\sqrt[3]{0.01}\] \[\sqrt{8}^{\log \frac{ 1 }{ 2 }}\] \[\frac{ 1 }{ \log _{_{8}}12-\log _{8} 3}\] \[(\log _{6}3) (\log _{3}36)\]